### v3

# STORM: Token-Efficient Long Video Understanding for Multimodal LLMs

Abstract
Recent advances in video-based multimodal large language models (Video-LLMs) have significantly improved video understanding by processing videos as sequences of image frames. However, many existing methods treat frames independently in the vision backbone, lacking explicit temporal modeling, which limits their ability to capture dynamic patterns and efficiently handle long videos. To address these limitations, we introduce STORM (Spatiotemporal TOken Reduction for Multimodal LLMs), a novel architecture incorporating a dedicated temporal encoder between the image encoder and the LLM. Our temporal encoder leverages the Mamba State Space Model to integrate temporal information into image tokens, generating enriched representations that preserve inter-frame dynamics across the entire video sequence. This enriched encoding not only enhances video reasoning capabilities but also enables effective token reduction strategies, including test-time sampling and training-based temporal and spatial pooling, substantially reducing computational demands on the LLM without sacrificing key temporal information. By integrating these techniques, our approach simultaneously reduces training and inference latency while improving performance, enabling efficient and robust video understanding over extended temporal contexts. Extensive evaluations show that STORM achieves state-of-the-art results across various long video understanding benchmarks (more than 5% improvement on MLVU and LongVideoBench) while reducing the computation costs by up to and the decoding latency by 2.4-2.9 for the fixed numbers of input frames. Project website is available at here.
1 Introduction
Recent advancements in video-based multimodal Large Language Models (Video-LLMs) have improved AI systems’ ability to understand video content [1, 2, 3, 4, 5, 6]. These models typically process videos as sequences of individual frames, encoding each frame independently before feeding representations to a large language model (LLM) for temporal reasoning [3, 4, 2]. Despite leveraging powerful image encoders and LLMs [7, 8], these approaches exhibit fundamental limitations in video processing, particularly for long videos. The lack of explicit temporal encoding in image tokens forces LLMs to infer temporal relationships from static image sequences, creating substantial computational burden. Moreover, to manage computational overhead, existing methods like LongVILA [9] and LongVA [10] often use naive frame subsampling to reduce LLM processing tokens, resulting in significant information loss, potentially eliminating critical details necessary for comprehensive video understanding. Recently, LongVU [11] proposes token compression by adaptively selecting visual tokens based on user queries. However, unlike our approach, LongVU lacks a dedicated module for propagating spatiotemporal information before token selection, potentially limiting the model’s ability to preserve critical information after token reduction.
In this paper, we propose STORM (Spatiotemporal TOken Reduction for Multimodal LLMs), which introduces a novel temporal encoder between the image encoder and the LLM. This design integrates temporal dynamics earlier in the model pipeline, significantly enhancing the temporal reasoning capabilities of Video-LLMs while enabling substantial downstream token reduction. We employ Mamba [12] as the core of our temporal layer to allow both computation efficiency and generalization to extended temporal contexts. While there are existing Mamba-based video understanding models [13, 14, 15, 16, 17], unlike ours, these methods do not integrate Mamba with LLMs for multimodal tasks, and importantly, they mainly focus on replacing traditional backbones with Mamba architectures without explicitly leveraging Mamba’s ability for reducing video redundancy and enabling visual token compression.
The key advantage of the Mamba layer is its ability to compress historical information into state representations. As consecutive frames in the video input often contain redundant information, our temporal encoder efficiently processes and propagates temporal information across the entire video sequence. The resulting visual tokens inherently encapsulate temporal history, allowing us to extract fewer tokens for LLM processing while preserving key information. We explore both training-free token subsampling and training-based compression methods (temporal and spatial token compression). Unlike previous frame-subsampling approaches, our method preserves essential temporal information in a compressed format, reducing computational load while improving performance through more comprehensive representation in a compact token space. Empirical evaluations demonstrate that our method significantly outperform baselines, especially on long-video inputs, while simultaneously achieving strong computational efficiency.
2 Related Work
Vision Language Model Architecture
Vision language models (VLMs) primarily adopt two paradigms to process visual input. The first paradigm freezes language model weights and integrates visual information via cross-attention mechanisms [18], while the second paradigm utilizes a pre-trained image encoder, such as CLIP [7] or SigLIP [8], to convert images into tokens. These tokens are then concatenated with text tokens and input into the language model [2, 1, 19]. This approach can be naturally extended to video understanding by treating videos as sequences of images processed by the vision encoder [3, 4]. To enhance video processing, some works introduce specialized video encoders. For instance, InternVideo [5, 6] uses VideoMAE [20] as a video encoder, while Kangaroo [21] integrates depth-wise 3D convolution for fusing video tokens. In this work, we retain SigLIP as the vision encoder and focus on enhancing long video understanding by incorporating a linear-complexity temporal module in the Mamba [12] architecture. Positioned between the SigLIP vision encoder and the language model, this module efficiently improves spatial-temporal modeling effectiveness.
Long Video Understanding
Understanding long videos with VLMs presents significant challenges in both accuracy and efficiency. Previous approaches have employed long-context language models trained on short-context video data to enable long video comprehension [10]. However, these methods lack sufficient long video training data and incur high computational costs during both training and inference as the number of frames increases. LongVILA [9] addresses these challenges through a multi-modal sequence parallelism system that directly handles long video data during training and inference, but this approach requires customized system implementations tailored for multi-GPU setups. Another line of research focuses on token reduction to shorten input sequences, thereby enabling efficient inference for long videos [22, 23, 24, 25, 26, 11]. For instance, VoCO-LLaMA [24] and VideoXL [26] use recursive KV cache compression learnt in an end-to-end manner, and LongVU [11] leverages DINO features for frame selection and inter-frame similarity to reduce tokens. Despite these diverse strategies, direct pooling along the temporal or spatial dimensions often performs sufficiently well, with additional gains being marginal. In this paper, we apply temporal and spatial pooling for token reduction, achieving superior performance when combined with our temporal projector.
Mamba for Video Understanding
Recent advances in linear state space models such as Mamba [12, 27] have sparked extensive exploration in applying them to video understanding tasks. Due to their sub-quadratic computation complexity, Mamba models achieve significant efficiency improvements compared to transformer-based architectures while still delivering competitive performance [12, 27, 28, 29, 30]. These properties make Mamba particularly suitable for video processing as the models are required to process long sequence inputs. For example, VideoMamba in [13] and VideoMamba (identical model naming) in [14] use Mamba-based visual backbone in video models and demonstrates the model’s strong ability to capture both local redundancy and long-term spatiotemporal dependencies. VideoMambaPro [15] proposes to improve Mamba’s video understanding ability by applying masking and residual connection during the backward scan. The Video Mamba Suite [16] further explores various architectures to integrate Mamba into existing video understanding models, demonstrating favorable efficiency-performance trade-offs for long sequence inputs. Mamba-ND [17] aims to improve Mamba’s performance on multi-dimensional data by investigating design choices such as SSM layer structure and scanning order within and across dimensions. However, unlike our approach, these works do not directly apply Mamba for Multimodal LLMs. More importantly, they primarily focus on replacing traditional backbones with Mamba architectures without explicitly leveraging Mamba’s unique ability to summarize historical information for reducing video redundancy and enabling visual token compression. Our paper addresses this research gap by proposing STORM, which proves to be both effective and efficient for video understanding while significantly reducing computational demands.
Concurrent Work
Recently, BIMBA [31] explores a similar architecture for long-video understanding, reporting similar benefits through empirical evaluation. We are encouraged to see these independent findings further support the hypothesis.
3 Method
This section introduces the Mamba-based temporal projector architecture and explores several token compression techniques for efficient long video processing. We investigate both training-based and training-free token compression strategies across temporal and spatial dimensions. We begin with an overview of our Mamba-based temporal projector, followed by a discussion of the token compression methods. The overview of our method is shown in Figure 2.
3.1 Preliminaries
State Space Models (SSMs) A State Space Model (SSM) [32, 33, 12, 34] establishes a linear transformation between an input sequence and an output sequence through the following recurrent process:
Here, is the sequence length; are input and output vectors at time ; and is the hidden state summarizing the history . The matrices , , and are parameterized with learnable weights designed to facilitate the modeling of long-term dependencies. When are time-invariant (constant over ), the computation of can be parallelized, enabling efficient training and inference.
Mamba Recently, Mamba [12] propose to condition these matrices on the input to enhance the sequence modeling capabilities of SSMs. Specifically, Mamba employs learnable functions , , and to generate input-dependent matrices as follows:
This approach allows the model to dynamically emphasize or suppress information based on the current input, thereby enabling more flexible and adaptive sequence modeling. Additionally, Mamba leverages a hardware-aware parallel algorithm to ensuring that the input-dependent matrices does not hinder the training and inference efficiency inherent to SSMs.
3.2 Mamba-Based Temporal Projector
Traditional Video-LLMs often process video frames independently, requiring the LLM to infer temporal relationships from sequences of static images. This approach is computationally inefficient, particularly when processing long videos. Additionally, this method fails to leverage the inherent temporal redundancy in video data, resulting in redundant processing of similar information across consecutive frames. To address these limitations, we introduce the Mamba-based temporal projector, which efficiently integrates temporal information across video frames while enabling effective temporal token compression.
Let denote the image tokens for frame by a Vision Transformer (ViT) encoder, where is the number of tokens per frame and is the token dimension. We first apply a linear layer to downsample the tokens of each frame to tokens:
Where is the downsample ratio. For simplicity, we define and use throughout the remainder of this paper. The downsampled tokens from all frames are stacked to form the input tensor for the temporal module:
The core of the temporal projector consists of Mamba layers that iteratively integrate temporal dynamics into the tokens. In each layer , we fuse temporal information into the visual tokens by:
where , and denotes layer normalization. Each MambaMixer employs a bidirectional scanning module that captures dependencies in both spatial and temporal dimensions. Specifically, we apply a sweeping scan order within each frame and across frames, i.e., left-to-right, top-to-bottom, and frame-to-frame (see Figure 2). After layers, we obtain tokens enriched with temporal information, denoted as .
3.3 Training-Time Token Compression
Processing long videos poses two major challenges for LLMs. First, handling all frames is computationally expensive, often requiring specialized system optimizations like sequence parallelization and multiple GPUs for training and inference [9]. Second, LLMs are inherently limited by their training context lengths constraints. For example, LLaMA 3 has a context length of 8K tokens, which corresponds to only 32 frames for video inputs when using 256 tokens per frame. Without token compression, video processing quickly exceeds the effective input capacity of LLMs, leading to significantly reduced model performance. In this work, we aim to address both computational cost and context length limitations by enabling efficient long-video processing through token compression along both temporal and spatial dimensions. Our approach eliminates the need for custom system optimizations and allows inference on a single GPU. We illustrate the training-time token compression in Figure 3 (left and middle).
Temporal Pooling
Since consecutive frames often contain similar information, analyzing every frame may lead to redundant processing and potential overfitting. Additionally, having too many tokens can make it difficult for the LLM to identify important temporal patterns. Thus, we propose applying temporal average pooling to compress visual information efficiently [35, 36]. This method combines data from consecutive frames by averaging their enriched visual tokens. Specifically, for the tokens from the temporal projector, we average every consecutive frames:
As a result, we obtain compressed tokens:
Despite its simplicity, temporal averaging effectively decreases the number of tokens the LLM processes, with minimal loss of critical information. This motivates us to adopt such simple yet effective technique for efficient training in long video understanding.
Spatial Pooling
In addition to temporal pooling, we also explore average pooling in the spatial domain. Formally, given the input from the visual encoder and the spatial compression ratio , we apply average pooling with a kernel size and stride of on each frame, resulting in .
3.4 Training-Free Temporal Token Sampling
After processing through the temporal projector, each visual token is enriched with spatiotemporal information, capturing features not only from its corresponding frame but also from other frames across the entire video. This encoding of global information allows us to subsample the visual tokens along the temporal dimension at test time, further reducing the number of tokens fed into the LLM without significant loss of information or performance. This temporal token subsampling strategy can be done with or without our pooling mechanisms. Note that comparing to methods that subsample raw frames and risk discarding critical temporal cues, our approach leverages the spatiotemporal richness of tokens post-temporal encoding. We show an illustration of the temporal token sampling in Figure 3 (right).
Formally, let denote the token input to the subsampling layer which can be tokens from visual encoder , or tokens output from compression modules or . Here, and are the temporal and spatial dimensions after any temporal or spatial pooling. We apply a uniform subsampling with rate along the temporal dimension.
Our empirical results demonstrate that this subsampling method not only maintains performance on various video understanding benchmarks but can also improve it by reducing noise from redundant frames.
4 Experiments
In this section, we extensively evaluate the proposed method on various video understanding benchmarks and provide empirical analysis demonstrating how the temporal projector enables efficient token reduction while delivering strong video reasoning abilities.
4.1 Experiment Details
We use the pre-trained vision encoder in PaliGemma [37] and the LLM model from Qwen2-VL [38], and fine-tune them to adapt to our video datasets. The temporal projector is initialized with random weights. Each image is always resized to a resolution. Note that we also carried out experiments on more model configurations in Section 5 to showcase the universality of our design.
Models Size Comp. #Frames # Frames MVBench MLVU LongVideoBench VideoMME Ratio (%) (train) (test) test dev val (w/o sub.) Duration 16 sec 3120 min 8 sec60 min 160 min Proprietary Models GPT4-V - - 1fps 43.7 - - 60.7 GPT4-O - - 1fps 64.6 66.2 - 77.2 Open-Source Video-LLMs Video-LLaVA [39] 7B - 8 8 41.0 47.3 - 40.4 LLaMA-VID [22] 7B - 1fps 1fps 41.9 33.2 - - Chat-UniVi [40] 7B - 64 64 - - - 45.9 ShareGPT4Video [41] 8B - 16 16 51.2 46.4 - 43.6 LLaVA-NeXT-Video [42] 7B - 16 32 33.7 - - 46.5 VideoLLaMA2 [4] 7B - 16 32 54.6 48.5 - 46.6 VideoChat2 [3] 7B - 8 16 60.4 47.9 - 54.6 Long-LLaVA [43] A13B (53B) - - 128 54.6 - - 51.6 LongVA [10] 7B - - 128 - 56.3 - 54.3 LongVILA [9] 7B - 2048 256 - - - 57.5 mPLUG-Owl3 [44] 8B - 8 128 54.5 - 52.1 53.5 LLaVA-OneVision [45] 7B - 32 32 56.7 64.7 56.5 58.2 Kangaroo [21] 8B - 160 64 61.0 61.1 54.8 56.0 VideoXL [26] 7B - 128 2048 55.3 64.9 49.5 55.5 Oryx-1.5 [46] 7B - 256 64 67.6 67.5 56.3 58.8 LongVU [11] 7B - - 1fps 66.9 65.4 - 60.6 Qwen2-VL [38] 7B - 2fps 2fps 67.0 - 55.6 63.3 Ours 7B 25/12.5* 128 256 71.3 72.9* 60.5* 63.4 * test-time temporal sampling is applied, 2 additional compression.
Training
In the first stage, known as the Alignment Stage, we freeze both the image encoder and the LLM, training only the temporal projector using a small image-text dataset [47], containing 95K image-text pairs. Note that the Mamba layers perform not only temporal scan but also spatial scan within images, so video inputs are not strictly required to train it. For alignment stage, we find it sufficient to use only image-text pairs to pretrain the temporal projector. In the second stage, the supervised fine-tuning stage (SFT), we fine-tune all three components using a large and diverse dataset that includes text-only, image-text, and video-text data. There are around 12.5M samples in our SFT data mixture. due to space constraints. At this stage, we use 32 frames for each video input. For models with training-time token compression , we use a compression ratio of 4 — temporal pooling models compress 32 frames to 8 frames while spatial pooling models compress 256 tokens per image into 64 tokens. Moreover, for models with training-time token compression , we further employ a long video fine-tuning stage using 128-frames long-video inputs from the LLaVA-Video dataset [48]. We provide further details about the full SFT dataset and long video fine-tuning dataset in the appendix Section 8.4.
VILA-Based Models Size Comp. Latency #Frames # Frames MVBench MLVU LongVideoBench VideoMME Ratio (%) (s) (train) (test) test dev val (w/o sub.) Duration 16 sec 3120 min 8 sec60 min 160 min Token Budget: 8K VILA Baseline 7B 100 4.31 32 256 69.5 70.2 55.9 60.1 STORM 7B 100 4.47 32 256 70.3 71.1 54.5 62.4 STORM + T. Pooling 7B 25 1.82 128 256 71.3 72.5 59.5 63.4 STORM + T. Sampling * 7B 50 2.50 32 256 70.1 70.8 54.8 63.1 STORM + T. Pooling + T. Sampling * 7B 12.5 1.51 128 256 70.6 72.9 60.5 62.4 * 2 additional compression at test time.
Evaluation
We evaluate our STORM across multiple configurations on recent long-video understanding benchmarks specifically designed to rigorously assess video-language model capabilities. These benchmarks include MVBench [49], MLVU [50], LongVideoBench [51], and VideoMME [52]. We compare our approach against a wide range of representative video-language models, including recently proposed models tailored for long-video understanding [9, 11, 10, 26, 43, 21, 46]. See Table 1 for details.
Models
We implement STORM based on the VILA codebase [2], a typical VLM pipeline consisting of a vision encoder, LLMs, and vision-language projector, and introduce our novel Mamba module and compression mechanisms into the architecture. We will refer all models trained with the VILA codebase as VILA-based models in our experiments. To thoroughly analyze our design, we evaluate STORM variants using all combinations of the three compression methods: temporal average pooling, spatial average pooling, and temporal token sampling. The full results are presented in Table 5. For comparison with existing Video-LLMs, we highlight the best-performing variants in Table 1. And include the detailed analysis of all variants of STORM in Table 2. To ensure fairness, we also include a baseline VILA model trained on the same dataset and training scheme but without the Mamba module. All of our models in Table 2 are trained under the same 8K token budget (corresponding to 32 frames of tokens ), which represents the number of visual tokens fed to the LLM—a key factor affecting inference latency and memory consumption, particularly as the number of frames increases (see Section 5). Specifically, we report results for the standard STORM trained on 32-frame inputs and STORM with Temporal Pooling, which processes 128-frame inputs while reducing the token count to match the 32-frame variant. Additionally, we evaluate configurations where temporal sampling is applied at test time (+T. Sampling), which not only further enhances model efficiency but also improves performance on certain benchmarks.
4.2 Results on Video Understanding Benchmarks
STORM vs Existing Methods
We start with comparing our configuration STORM + T. Pooling (+ T. Sampling) with existing Video-LLMs. As shown in Table 1 and detailed in Table 2, STORM + T. Pooling achieves new state-of-the-art performance across all long-video understanding benchmarks. Specifically, it achieves 71.3% accuracy on MVBench, 72.5% on MLVU, 59.5% on LongVideoBench, and 63.4% on VideoMME, outperforming all open-source Video-LLMs, including recent models specifically designed for long-context inputs such as LongVU and LongVILA. Additionally, our method significantly narrows the performance gap with proprietary models, outperforming GPT4-V and GPT4-O on MVBench and MLVU, as well as GPT4-V on VideoMME. Notably, STORM + T. Pooling achieves computational efficiency by compressing visual tokens to 25% of their original number before processing them through the LLM. We can further enhance efficiency by applying test-time temporal sampling in STORM + T. Pooling + T. Sampling while still achieving competitive results, reducing the token count to just 12.5% while maintaining competitive performance. In fact, this additional compression even improves results of certain benchmarks, yielding the best overall performance on MLVU and LongVideoBench.
STORM vs Baseline VILA
Next, we provide controlled comparisons within VILA-based models to reveal advantages of the proposed Mamba-based temporal module in Table 2. We first compare the baseline VILA model with our STORM. By incorporating the Mamba module, STORM achieves performance gains on three out of four benchmarks, including a notable 2.3% improvement on VideoMME. Moreover, augmenting STORM with test-time temporal token sampling (STORM + T. Sampling) further enhances efficiency, reducing inference time by 43% while, surprisingly, maintaining or slightly boosting performance (an additional 0.8% gain on VideoMME). This advantageous behavior emerges because the Mamba module’s ability to effectively propagate temporal information across video frames, enabling redundant tokens to be discarded without compromising the model’s overall understanding.
The temporal pooling variant (STORM + T. Pooling) extends these benefits to long-context training, by applying temporal average pooling after the Mamba layer, which allows the model to process 128-frame inputs while compressing the token count to match that of the 32-frame setting. This approach not only improves performance, achieving 63.4% (+3.3%) on VideoMME, 59.5% (+3.6%) on LongVideoBench, 72.5% (+2.3%) on MLVU, and 71.3% (+1.8%) on MVBench, but also significantly reduces inference latency by 58.4%. By combining this model with test-time temporal token sampling (STORM + T. Pooling + T. Sampling), we further reduce the inference time by 65.5% and use only 12.5% of the visual tokens compared to VILA without sacrificing performance. We observe that this test-time temporal token sampling particularly benefits MLVU and LongVideoBench compared to MVBench and VideoMME. This different impact across benchmarks likely stems from the nature of the underlying tasks. MLVU and LongVideoBench require global understanding across long videos. We believe that test-time compression with the Mamba module better summarizes the essential contextual information. On the other hand, MVBench and VideoMME require visual details from specific frames. Our pooling-only method with the Mamba module maintains more detailed frame information throughout the sequence. Section 5 includes a detailed discussion about retaining visual information with our token compression.
Spatial Pooling vs Temporal Pooling on Long Video Inputs
Table 3 provide comparison between spatial average pooling and temporal average pooling on 32 frames training and 128-frames extension. All models use STORM as the base model. We see that spatial pooling is effective when 32 frames are during training, outperforming temporal pooling and on LongVideoBench while achieving on par results on VideoMME. However, when applying 128-frames input, even though both methods use the same token budget, temporal pooling results in sigificantly better performance. In fact, spatial pooling can not benefit from longer video inputs, results in degrated performance on both LongVideoBench and VideoMME, while temporal pooling successfully achieves stronger performance on both benchmarks from the extended video length.
Models LongVideoBench VideoMME 32F 128F 32F 128F S. Pooling 56.0 55.9(-0.1) 61.1 58.3 (-2.8) T. Pooling 54.2 59.5 (+5.3) 61.2 63.4 (+2.2)
5 Analysis
Models 8F 32F (T. Pooling) 128F (T. Pooling) VILA (w/o Mamba) 62.0 63.5 (+1.5) 64.3 (+0.8) STORM (w/ Mamba) 61.6 64.2 (+2.6) 66.7 (+2.5)
Mamba Module Improves Simple Token Compression
As illustrated in Figure 2, the Mamba-based temporal projector performs a spatiotemporal scan over input tokens. This enables refinement of visual tokens before compression operations, thereby preserving key temporal and spatial cues—such as the positional information of frames being pooled—that would otherwise be lost in naive compression strategies. Consequently, the Mamba temporal module allows simple token compression methods to work more effectively for long videos compared to the baseline, as evidenced in Table 4 and further detailed in Table 10.
Our results demonstrate that STORM consistently improves with increased video input length during training, achieving substantial gains of +2.6% over all benchmarks when extending from 8 frames to 32 frames and an additional +2.5% when extending from 32 frames to 128 frames. In contrast, the baseline VILA model shows smaller improvement when extending from 8 frames to 32 frames, and demonstrates only +0.8% performance gains when fine-tuned on 128 frames. In fact, as shown in Table 10, the baseline model exhibits performance degradation on MVBench and MLVU benchmarks when extending video length from 32 frames to 128 frames. These results fully demonstrate the importance of the Mamba module in enabling both efficient and effective token compression, particularly for long-video input
Model Mamba T. T. S. Token Compression Latency (s) MVBench MLVU LongVideoBench VideoMME Module Sampling Pooling Pooling budget Ratio (%) w/o sub. VILA ✗ ✗ ✗ ✗ 8K 100 4.31 69.5 70.2 55.9 60.1 STORM ✓ ✗ ✗ ✗ 8K 100 4.47 70.3 71.1 54.5 62.4 ✓ ✓ ✗ ✗ 8K 50 2.50 70.1 70.8 54.8 63.1 ✓ ✗ ✓ ✗ 2K 25 1.82 70.4 71.0 54.2 61.2 ✓ ✗ ✗ ✓ 2K 25 1.82 68.9 69.2 56.0 61.1 ✓ ✓ ✓ ✗ 2K 12.5 1.51 70.1 71.0 54.0 60.8 ✓ ✓ ✗ ✓ 2K 12.5 1.51 68.9 69.5 56.3 60.9 ✓ ✗ ✓ ✓ 0.5K 6.25 1.36 68.5 68.2 53.7 60.2 ✓ ✓ ✓ ✓ 0.5K 3.13 1.29 67.2 68.6 55.3 60.1
Ablation of Mamba and token compression modules
Table 5 presents an ablation study comparing various token compression strategies for models trained on a fixed 32-frame input. Without any compression, our standard STORM delivers improved performance across benchmarks comparing to the baseline VILA while utilizing a similar inference latency and number of visual tokens. When test-time temporal sampling is applied, the model maintains its performance while reducing the visual token count to 50% and lowering inference latency to 58.7% of the uncompressed STORM. Employing temporal average pooling during both training and testing further compresses the token count to 25% and cuts latency down to 42.7%, which is particularly effective for MVBench and MLVU but slightly degrades LongVideoBench and VideoMME. Similarly, spatial average pooling provides the same efficiency improvements and proves particularly effective on LongVideoBench, but with compromises on other benchmarks.
We further apply temporal token sampling with either temporal or spatial pooling to improve the test-time efficiency. For instance, STORM + T. Pooling + T. Sampling reduces the visual tokens to only 12.5% and latency to 35.4% of the original STORM, while maintaining comparable accuracy to STORM + T. Pooling. The STORM + S. Pooling + T. Sampling even achieves the best LongVideoBench performance with this strong token reduction. Finally, we combine the temporal and spatial average pooling with/without test-time temporal token sampling, leading to even more aggressive compression. Interestingly, our STORM + T. Pooling + S. Pooling + T. Sampling variant, which has the strongest compression ratio, utilizing only 3.13% visual token and 29.5% of inference latency, already achieves a competitive results. Comparing to the baseline VILA, it achieves 100% performance on VideoMME, 98.9% on LongVideoBench, 97.7% on MLVU, and 96.7% on MVBench, making it quite attractive for efficiency-oriented scenarios.
We note that, although the uncompressed STORM and its test-time sampling variant achieve the highest overall performance when using the same 32 frames input, their computational demands limit scalability to train on longer sequence. In contrast, as demonstrated in Table 2, the STORM + T. Pooling allows extending the temporal context to 128 frames with temporal pooling, which enables improved performance without requiring additional LLM computation.
Token Compression Improves Model Efficiency
We provide analysis on model efficiency in Figure 4. In Figure 4 left, we compare the inference latency of STORM with and without token compression and across varying numbers of input frames. Due to the quadratic computation of the LLM, the latency gap between the two configurations widens significantly as the sequence length increases. Figure 4 middle provides a detailed breakdown of inference latency of different modules for processing 256 frames. Without token compression, approximately 80% of the total latency is attributed to the LLM module. By applying a token compression ratio of 4, the latency associated with the LLM is reduced to roughly 35% of its original value, demonstrating a substantial improvement in efficiency.
Importantly, the temporal projector enables token compression without compromising performance. In fact, for some benchmarks, it even improves model performance while reducing computation cost, as demonstrated in Table 2 and detailed in Table 9. Figure 4 right presents a detailed analysis using the temporal sampling method on the VideoMME benchmark. We compare three configurations: the VILA baseline, STORM without compression, and STORM with test-time temporal token sampling Using up to 128 input frames and a compression ratio of 2. The results indicate that: (1) The VILA baseline’s performance improves up to 32 frames but declines beyond this point, (2) STORM generalizes better to longer sequences, maintaining performance up to 64 frames before a slight drop, and (3) STORM with temporal sampling continues to improve performance up to 128 frames while providing a 50% token compression ratio, meaning STORM + T. Sampling at 128 frames uses the same number of visual tokens as the other two models at 64 frames.
We speculate that phenomenon is because of the limitation of the pre-trained LLM models which, despite having large context lengths on paper, operate with a smaller effective context length in practice. This limitation hinders their performance on very long sequences. The combination of the temporal projector and temporal sampling ensures that the number of tokens fed into the LLM remains within its effective context length. Simultaneously, the enriched visual tokens retain comprehensive temporal information beyond the sampled frames.
Mamba Module Has Minimal Training Overhead
The Mamba projector introduces minimal training overhead: even when no compression is applied, it adds only 5 latency in full training (19.1 hours with VILA (without Mamba module) and 20.1 hours with STORM (with Mamba module)). More importantly, since the Mamba module enables effective token compression, it in fact provides significant training / inference speedups and performance gains (e.g., +3.6 on MLVU with 4× fewer tokens), eventually reducing training the costs.
Our Token Compression Retains Visual Information
The qualitative results in Figure 5 demonstrate that STORM’s token compression preserves critical visual information while significantly reducing computational overhead. Even with 4 compression ratio, our model accurately extracts task-relevant information across diverse video understanding tasks. For example, questions like ’Whom is the poem in the video written by?’, ’How many times do news segments appear in this video?’, and ’What is unique about the last performance?’ require fine-grained visual information and focus on the specific content from the prompt. Detailed category-level analysis on VideoMME (Figure 7) also reveals that STORM + token compression consistently outperforms baseline model across nearly all task categories. Even in OCR-heavy tasks that require detailed visual analysis, our compressed model maintains performance comparable to uncompressed versions while using only a fraction of the computational resources (only 25% visual tokens). These findings demonstrate that our Mamba-based temporal encoder enables effective token compression by encoding spatiotemporal relationships directly into visual tokens, allowing the model to maintain high-level understanding while drastically reducing token count. Additional qualitative results are provided in Figure 10-13.
Our Modules Generalize to Various Architectures
Our core design, temporal module + token compression, is model-agnostic and can be integrated into various video-LLM architectures. Table 6 shows the experiments on models utilizing different LLM models (Qwen2 [38] vs Llama3 [53]), vision encoders (PaliGemma [37] vs SigLip [8]), model sizes (8B vs 8B vs 1.5B), and input resolutions ( vs ). Our design shows consistent improvements in both performance and latency for all configurations, clearly showcasing the universality of the model.
Architectures Resolution Latency (ms) VideoMME VILA Ours VILA Ours Qwen2 7B + PaliGemma 448448 1980 926 59.7 61.2 Llama3 8B + SigLip 384384 1560 724 54.6 56.8 Qwen2 1.5B + SigLip 384384 724 486 49.6 52.6
Models 32F (T. Pooling) 128F (T. Pooling) VILA 58.9 61.7 Uni-dir STORM 62.2 62.5 Bi-dir STORM 61.2 63.4
Support for Streaming/Online Settings.
To evaluate the applicability of our method in streaming scenarios, we replaced the default bi-directional Mamba with a uni-directional variant, allowing the model to reuse prior states for constant-time computation as new frames arrive. The results are provided in Table 7. We find that both uni-directional and bi-directional variants significantly outperform the baseline without temporal modeling. Notably, the uni-directional Mamba performs competitively and even slightly surpasses the bi-directional counterpart when trained on 32-frame inputs. On the other hand, the bi-directional model demonstrates stronger performance as video length increases. These results highlight the critical role of the Mamba module and suggest its potential in enabling future designs that support streaming video input for video-LLMs.
6 Conclusion
We introduced STORM, a novel Video-LLM model that enhances long-video understanding using a Mamba-based temporal encoder and efficient token reduction. By explicitly integrating spatiotemporal dynamics into visual tokens early in the pipeline, STORM enables significant token compression while preserving critical information in the compressed inputs. Experiments demonstrate that STORM achieves new state-of-the-art results on long-video understanding benchmarks while substantially improving computational efficiency.
7 Acknowledgments
We would like to thank Xin Dong, Matthieu Le, Yunhao Fang, and Dacheng Li for their valuable discussions on this work. Sungjin Ahn was supported by the NRF of Korea funded by the Ministry of Science and ICT (MSIT) (No. 2021H1D3A2A03103645).